# PIP-Net: Pedestrian Intention Prediction in the Wild

### v2

@@ IV-C2 Features Importance
Excluding Semantic Context leads to a 3.4% accuracy decrease as it includes details about road layout such as sidewalk positioning and drivable zones. The impact of removing Local Content is minor, causing accuracy to decrease by 1.7%. It appears to lack comprehensive features on pedestrians’ intentions, given the wide variety in appearance and accessories they may have.
Table III shows a comparison of the two input features we have used in the proposed model, which correspond to the scene motion and depth information. The $\alpha^{2}$ demonstrates that the local motion feature can exhibit superior performance when compared to the global motion feature (2d), which includes optical flow analysis of the entire scene. While optical flow is typically sensitive to any movement between consecutive frames, it appears that local motion can provide a coarse-grained feature and treat more concisely to account pedestrians’ velocity and direction of movement, regardless of other irrelevant objects in the surroundings.
Regarding depth information features, the categorical depth feature proposed stands out as the most effective standalone feature, as evidenced by the results of $\alpha^{4}$ in Table III, highlighting the importance of pedestrian group densities, their distances from the ego-vehicle, and interactions with other road users in the traffic scene. Conversely, $\alpha^{3}$ performs the weakest among the sub-variants when compared with the baseline, which utilises the global depth heatmap of the entire scene (2e). This underperformance may be attributed to the unstable estimation of depth for irrelevant surrounding objects. This is addressed in the categorical depth by focusing only on pedestrians and vehicles, and then applying per-instance normalisation (as shown in Figure 4).
Finally, the optimal outcome is attained by taking into account both the local motion and categorical depth map, leading to enhancements in the baseline regarding Acc, AUC, and F1 score by 1.8%, 3.1%, and 6.1%, respectively.
IV-C3 Generalisation
The PIP-Net-$\alpha$ evaluation is presented against the SOTA methodologies on the Frontal-Urban-PIP dataset in Table IV. Notably, the models have never seen the scenarios in their training phase. The outcomes of other methods were generated using the pre-trained weights they provided. Overall, the majority of models demonstrate improvements over ATGC across various metrics, with each model exhibiting its unique strengths. However, we witnessed a decrement in terms of performance for PCPA, PPCI, CAPformer, and GraphPlus. As far as the research curiosity demand, we investigated the architecture of these models, and it turned out they are suffering from low-quality global context and body pose features. This is caused by the feature extractor algorithms, i.e., semantic segmentation and pose estimator, they have used which hinder the classifier from judging based on precise features.
IV-C4 View Angle Expansion
@@
We explored the enhancement of the field of view using three cameras to enable the autonomous vehicle to perceive a larger portion of its surroundings. For this purpose, we train PIP-Net-$\beta$ with three different observation times ($m$) of 20, 30, and 40 frames. Subsequently, we examined how the prediction performance evolves as we expand the ETC prediction horizon from 1 to 4 seconds. As depicted in Figure 7, the accuracy of crossing intention prediction decreases as the ETC prediction expands. However, the accuracy often increases with the enlargement of observation time. Intriguingly, when ETC = 4 and $m=40$ the accuracy was lower compared to $m=30$. This discrepancy arises from the model predicting pedestrians to be crossing based on long-term observations when, in reality, they did not cross. This highlights the observation that a pedestrian’s previous actions do not always accurately indicate their future intentions, as they can change their mind and act in an instant [3].
Figure 7: Comparison of the crossing intention accuracy (Acc) for PIP-Net-$\beta$ models (in three cameras mode) across various future time intervals, highlighting the disparities in performance at each observation time ($m$). The results are based on the Urban-PIP dataset.
IV-D Observational Results
@@
We present the qualitative results of the PIP-Net-$\alpha$ network in Figure 8 for Frontal-Urban-PIP datasets and in Figure 9 for the PIE dataset. The intention is represented by a confidence bar, where higher values (reddish colour) indicate a high probability of the pedestrian crossing. Pedestrians without the intention to cross are depicted with greenish bounding boxes and lower values on the confidence bar.
@@
(b) A signalised junction with a green light
@@
Figure 8: A representation of the proposed model’s analysis on traffic video scenes of Frontal-Urban-PIP dataset. Three distinct images showcase intersections and traffic light conditions with confidence bars positioned in front of pedestrians, aligning with their body pose direction. These confidence bars visually indicate the model’s confidence about the crossing intention of the detected pedestrians for the upcoming 2-4 seconds into the future.
We displayed frames from the past $t-90$ frames to the $t$ frame, where $t$ represents the decisive moment. As the frames change, the prediction results for the pedestrian’s crossing intention also evolve. Some pedestrians are predicted to continue crossing, while others are forecasted to transition from not crossing to crossing, or vice versa, depending on the pedestrian’s direction or situation. For instance, the proposed model can predict the intention of the pedestrian in case 7, as shown in Figure 9, when the ego vehicle is going to turn left and the pedestrian tends to cross. The contextual information alongside pedestrian features like motions and direction of movements, seems to be crucial for accurately predicting their behaviour in a given context.
Figure 9: Qualitative results of PIP-Net-$\alpha,(m=30)$ on the PIE dataset. The pedestrian with a crossing intention has shown a higher value in the confidence bar. We demonstrate the past frame $t-90\thicksim t$ and the ground truth as a time in the future ($t+\epsilon$) with dashed bounding boxes, where red: cross, green: not cross.
V Conclusion
@@
This paper presented a framework called PIP-Net for predicting pedestrian crossing intentions in real-world urban self-driving situations. Two variants, $\alpha$ and $\beta$, were introduced to support different camera setups. By utilising both kinematic data and spatial features of the driving scene, the proposed model employed recurrent and temporal attention-based methodology to predict pedestrians’ future crossing intentions accurately. Through quantitative and qualitative experiments on the PIE dataset, the proposed model achieved state-of-the-art performance with a 91% accuracy and an 84% recall rate.
@@
The Urban-PIP was introduced as a new dataset for the pedestrian intention prediction task, including various AV driving scenarios and comprehensive annotations on a multi-sensory setup, thereby enabling a better future investigation of crossing behaviour studies. Our model demonstrated a generalisation capability when applied to the Urban-PIP dataset by +9%, +10%, and 12% improvement compared with other models in terms of accuracy, AUC, and F1 score, respectively. This was underlined by the scene feature extractors employed in training our model.
@@
To enhance the visual encoding of road users and their relative distances to the ego vehicle, we introduced a categorical depth feature map. This, combined with the local motion flow feature, provided salient information about the dynamics of the scene. Our results reveal that they cumulatively enhanced the accuracy and F1 score of the baseline model by +2.8% and +5.2%, respectively. Additionally, we investigated the impact of expanding the view angle using three cameras and enlarging prior observation frames.
@@
We anticipate that this algorithm can effectively prevent traffic accidents and protect vulnerable road users by foreseeing the crossing behaviour of nearby pedestrians.

### v3

@@ IV-D3 Generalisation
Figure 8: Comparison of semantic and panoptic segmentation, where panoptic segmentation offers more precise scenes in daylight by distinguishing object instances and semantic categories, leading to improved global context representation.
Figure 9: Body pose estimation on local image rather than on the entire image enables more precise and reliable pose extraction.
We observed more precise body pose information by applying pose estimation on the local image region containing the pedestrian rather than the entire scene image. This method ensures a more focused and accurate pose representation, reducing noise and ambiguity caused by irrelevant background elements. Figure 9 illustrates how our approach successfully retrieves clear and precise pose data in scenarios where applying pose estimation on the entire scene fails to obtain reliable pose information. Applying the pose estimation algorithm to the local image region results in more robust and consistent pose features.
While these enhancements contribute to better generalisation, performance variations in challenging conditions emphasise the need for further research into feature extraction under adverse lighting and occlusions to ensure model reliability across all operational scenarios.
IV-D4 View Angle Expansion
@@
We explored the enhancement of the field of view using three cameras to enable the autonomous vehicle to perceive a larger portion of its surroundings. For this purpose, we train PIP-Net-$\beta$ with three different observation times ($m$) of 10, 20, and 30 frames. Subsequently, we examined how the prediction performance evolves as expand the ETC prediction horizon from 1 to 4 seconds. As depicted in Figure 10, the accuracy of crossing intention prediction decreases as the ETC prediction expands. However, the accuracy often increases with the enlargement of observation time. Intriguingly, when ETC = 4 and $m=30$ the accuracy was lower compared to $m=20$. This discrepancy arises from the model predicting pedestrians to be crossing based on long-term observations when, in reality, they did not cross. This highlights the observation that a pedestrian’s previous actions do not always accurately indicate their future intentions, as they can change their mind and act in an instant [3].
IV-D5 Computational Cost of Multi-Camera Integration
To assess the computational cost of integrating three cameras, we evaluated the inference time, memory usage, and parameter count of PIP-Net-$\beta$ with stitched multi-camera inputs. Our results indicate that while computational demands increase, the model remains suitable for real-time execution. Specifically, the three-camera version required 9.9M parameters, 2.9GB of memory, and an inference time of 127ms ($\approx$ 7.8 FPS). This represents an approximate 1.8× increase in model size, 2.1× in memory usage, and 2.3× in inference time compared to the single-camera setup (PIP-Net-$\alpha$). The feature integration module efficiently mitigates redundancy, ensuring that multi-camera integration remains feasible for real-world deployment.
IV-E Observational Results
@@
We present the qualitative results of the PIP-Net-$\alpha$ in Figure 11 for FU-PIP datasets. The crossing probability is represented by a probability bar, where higher values (reddish colour) indicate a high probability of the pedestrian crossing. Pedestrians with no intention to cross are depicted with greenish bounding boxes and lower values on the probability bar.
Figure 10: Comparison of the crossing intention accuracy (Acc) for PIP-Net-$\beta$ models (in three cameras mode) across various observation sequence lengths ($m$). The results are based on the Urban-PIP dataset.
@@
(b) A signalised junction with a blue light
@@
Figure 11: A representation of the proposed model’s analysis on traffic video scenes of FU-PIP dataset. Three distinct images showcase intersections and traffic light conditions with probability bars positioned in front of pedestrians, aligning with their body pose direction. These bars visually represent the model’s predicted probability of each detected pedestrian crossing within the next 2 to 4 seconds into the future.
Figure 12 presents the qualitative results of PIP-Net-$\alpha$ on 12 sample scenarios in the PIE and FU-PIP dataset. The model successfully predicted cases in rows (a) to (c), demonstrating robustness across varying daylight ambient lighting conditions. In row (c) of FU-PIP, the model accurately inferred the crossing intention of a pedestrian despite partial occlusion. However, in row (d) of PIE, the model struggled to make a correct prediction, when the pedestrian was nearly fully occluded.
In row (d) of FU-PIP, it seemed the model correctly predicted the intention of the pedestrian who stepped onto the road for crossing. However, the pedestrian subsequently abandoned the crossing and stepped back. In such rare cases, where the pedestrian ultimately does not cross, the ground truth is labelled as “not crossing”, consistent with the annotation criteria of the PIE and JAAD datasets.
In row (e) of PIE, the ego vehicle was accelerating, and a hesitation in crossing was observed in the pedestrian’s behaviour. Interestingly, the model predicted a 43% probability of crossing, but this is below the 50% probability threshold to be accepted as a crossing event. However, we hypothesised that the acceleration of the ego vehicle might have a dominant effect over visual observations. To investigate this, we applied the permutation feature importance approach, as suggested in [51], and replaced the speed value with a scenario in which the ego vehicle was decelerating. Notably, the model then predicted an 88% probability of crossing intention for this case. This finding aligns with previous studies suggesting that, in certain scenarios, the model prediction heavily relies on ego-vehicle speed.
In row (f) of PIE, the ego vehicle was preparing to turn left while the pedestrian was inclined to cross the intersecting road. This suggests that incorporating additional motion-related information about the ego vehicle, such as the heading angle/steering angle, could help to address such cases.
We also evaluated the model’s performance in nighttime scenarios, as shown in rows (e) and (f) of FU-PIP. In these cases, the model failed to classify pedestrian crossing intentions accurately. This poor performance can be attributed to two primary factors: (1) the reduced accuracy of input features extracted from panoptic segmentation and depth estimation under low-light conditions, and (2) the limited number of nighttime samples in the training dataset, which likely led to model underfitting in such scenarios.
Figure 12: Qualitative results on 12 sample scenarios. Rows with blue ticks represent successful predictions by the model and red crosses indicate wrong predictions. $t$ demonstrates the prediction at the decisive moment and ($t+\epsilon$) shows the ground truth (dashed bounding boxes). Red bounding boxes denote crossing, while blue denotes not crossing.
V Conclusion and Future Work
@@
This paper presented a framework called PIP-Net for predicting pedestrian crossing intentions in real-world urban self-driving situations. Two variants, $\alpha$ and $\beta$, were introduced to support different camera setups. By utilising both kinematic data and spatial features of the driving scene, the proposed model employed a recurrent and temporal attention-based methodology to predict pedestrians’ future crossing intentions accurately. Through quantitative and qualitative experiments on the PIE dataset, the proposed model achieved state-of-the-art performance with a 94% AUC and an 88% F1 score.
@@
The Urban-PIP was introduced as a new dataset for the pedestrian crossing prediction task, including various AV driving scenarios and comprehensive annotations on a multi-sensory setup, thereby enabling a better future investigation of crossing behaviour studies. Our model demonstrated a generalisation capability when applied to the Urban-PIP dataset by +9%, +10%, and 12% improvement compared with other models in terms of accuracy, AUC, and F1 score, respectively. This was underlined by the scene feature extractors employed in training our model.
@@
To enhance the visual encoding of road users and their relative distances to the ego vehicle, we introduced a categorical depth feature map. This, combined with the sliced motion flow feature, provided salient information about the dynamics of the scene. Our results reveal that they cumulatively enhanced the accuracy and F1 score of the baseline model by +3.1% and +6.2%, respectively. Additionally, we investigated the impact of expanding the view angle using three cameras and enlarging prior observation frames.
@@
Overall, this study paved the way for developing multi-modal solutions for pedestrian crossing prediction and provided insights to effectively protect vulnerable road users by foreseeing the crossing behaviour of nearby pedestrians. Following the approach taken by seminal works in the field this study also focused on RGB cameras; however, we extended the use of a single camera to multiple cameras (left, front, and right) to evaluate the effectiveness of contextual information expansion. As one of the possible future work, we recognise the potential benefits of integrating data from other sensors, such as LiDAR and radar, to further enhance the model robustness in adverse weather or challenging lighting conditions. Specifically, LiDAR can offer more precise and reliable depth estimation, enhancing the Categorical Depth representation beyond what can be achieved with monocular-based depth estimation algorithms.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgements
The authors would like to thank all partners within the Hi-Drive project for their cooperation and valuable contribution. This research has received funding from the European Union’s Horizon 2020 research and innovation programme, under grant Agreement No 101006664. The article reflects only the author’s view based on limited datasets and experiments, and neither the European Commission nor CINEA is responsible for any real-world use that may be made of the information this document contains.